Let fn(x)= x^n is not uniformly continuous on [0,1] but is uniformly continuous on [0,k]

1
Expert's answer
2022-03-14T17:51:21-0400

Solution


"{f_n}\left( x \right) = {x^n}" and "[0,1]"


For "x=1"


"{f_n}\left( 1 \right) = {1^n} = 1"


Therefore,


"\mathop {\lim }\limits_{n \to \infty } {f_n}\left( 1 \right) = \mathop {\lim }\limits_{n \to \infty } \left( 1 \right) = 1\"


Now for "0 \le x < 1"


"\begin{array}{c}
\mathop {\lim }\limits_{n \to \infty } {f_n}\left( x \right) = \mathop {\lim }\limits_{n \to \infty } {x^n}\\
\mathop {\lim }\limits_{n \to \infty } {f_n}\left( x \right) = 0 & & x < 1
\end{array}"


Therefore, 


"\mathop {\lim }\limits_{n \to \infty } {f_n}\left( x \right) = \left\{ \begin{array}{l}
0 & 0 \le x < 1\\
1 & x = 1
\end{array} \right.\"


This shows that the "{f_n}\left( x \right) = {x^n}" is piecewise convergent over the interval "[0,1]"


Now for the interval "[0,k]"


We have


"\begin{array}{c}
\mathop {\lim }\limits_{n \to \infty } {f_n}\left( x \right) = \mathop {\lim }\limits_{n \to \infty } {x^n}\\
\mathop {\lim }\limits_{n \to \infty } {f_n}\left( x \right) = \left\{ \begin{array}{l}
0 & x < 1\\
\infty & x > 1
\end{array} \right.
\end{array}\"


Therefore, for any value of "k>1" , the "{f_n}\left( x \right) = {x^n}" is divergent. 


Hence

"{f_n}\left( x \right) = {x^n}" is piecewise convergent over the interval "[0,1]"  whereas not uniformlay continuously on the interval "[0, k]"